### v5

@@ Appendix D Glossary (Key Terms & Acronyms)
Toxicity. Harmful, offensive, or unsafe outputs produced by generative models, typically evaluated through specialized benchmarks.
Deepfakes / Media Integrity. Synthetic or manipulated media generated by GenAI that can mislead or cause societal harm; a central integrity and governance concern.
Governance Frameworks. Formal standards and policies (e.g., NIST AI RMF, EU AI Act, ISO/IEC 42001, OECD AI Principles) that regulate responsible development and deployment.
Key Performance Indicators (KPIs). Quantifiable measures that assess AI responsibility and compliance. For example, calibration can be measured by the Expected Calibration Error (ECE):
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\,\Big|\text{acc}(B_{m})-\text{conf}(B_{m})\Big|,$
where $B_{m}$ denotes prediction bins, $\text{acc}(B_{m})$ the empirical accuracy, and $\text{conf}(B_{m})$ the mean confidence.
Agentic AI. Generative systems with planning, memory, and tool-use capabilities that enable autonomous decision-making but also introduce new safety and governance challenges.

### v6

@@ B.1. Glossary (Key Terms & Acronyms)
Toxicity Term Harmful, offensive, or unsafe outputs; typically evaluated via specialized benchmarks.
Deepfake Term Synthetic media mimicking real persons/events (face/voice/image/video).
Provenance Term Traceability of data/model/content origin; includes watermarking and content credentials.
System-level failure Term Failures from tool use, planning, memory, or multi-agent interaction.
Governance crosswalk Term Mapping technical evidence (benchmarks/KPIs) to regulatory obligations.
Privacy Term Protecting sensitive data from exposure; operationalized via DP, FL, and other PETs.
DP Acronym Differential Privacy: formal guarantee bounding information leakage about individuals.
FL Acronym Federated Learning: training across clients without centralizing raw data.
PETs Acronym Privacy-Enhancing Technologies: methods to reduce/mitigate privacy risks.
Membership inference Term Attack inferring whether a record was in a model’s training data.
Data extraction Term Recovery of memorized training snippets from a model.
PII Acronym Personally Identifiable Information: data that identifies an individual.
Model card Term Standardized documentation of a model’s intended use, data, metrics, and limits.
Datasheet Term Documentation for datasets (provenance, licensing/consent, collection, intended use).
Post-market monitoring Term Ongoing evaluation and incident response after deployment.
Risk families (C1–C5) Term Rubric: C1 Bias, C2 Toxicity, C3 Security/Robustness, C4 Mis/Disinfo, C5 Deepfakes.
Risk families (C6–C10) Term Rubric: C6 Privacy, C7 System failure, C8 Red-team realism, C9 Metric validity/docs, C10 Governance alignment.
NIST AI RMF Acronym U.S. National Institute of Standards and Technology AI Risk Management Framework.
EU AI Act Acronym EU regulation defining risk tiers and obligations (e.g., transparency, monitoring).
ISO/IEC 42001 Acronym AI management system standard for governance and operational controls.
SAIF Acronym Secure AI Framework (Google): security practices for AI systems.
AI RMF Acronym Artificial Intelligence Risk Management Framework (shorthand for NIST AI RMF).
GDPR Acronym General Data Protection Regulation (EU).
HIPAA Acronym Health Insurance Portability and Accountability Act (USA).
FDA Acronym Food and Drug Administration (USA).
OCC Acronym Office of the Comptroller of the Currency (USA).
SEC Acronym Securities and Exchange Commission (USA).
PRISMA Acronym Preferred Reporting Items for Systematic Reviews and Meta-Analyses.
SOTA Acronym State-of-the-Art.
IHL Acronym International Humanitarian Law.
NIST Acronym National Institute of Standards and Technology.
NATO Acronym North Atlantic Treaty Organization.
UNESCO Acronym UN Educational, Scientific and Cultural Organization.
SAG-AFTRA Acronym Screen Actors Guild–American Federation of Television and Radio Artists.
WGA Acronym Writers Guild of America.
AIF360 Acronym IBM AI Fairness 360: fairness metrics and mitigation toolkit.
Fairlearn Acronym Microsoft toolkit for fairness assessment and mitigation.
Theme Citation Year Focus
Safety/Benchmarks (Liu et al., 2024) 2024 Multimodal benchmark to assess safety risks in vision–language models.
Safety/Benchmarks (Zhang et al., 2024) 2024 General safety benchmark for LLMs across risk categories.
Safety/Benchmarks (Gehman et al., 2020) 2020 Toxic degeneration measurement using curated prompts to stress test LMs.
Safety/Benchmarks (Li et al., 2024a) 2024 Hierarchical, comprehensive safety benchmark for language models.
Safety/Benchmarks (Mazeika et al., 2024) 2024 Standardized evaluation for red teaming and refusal robustness.
Safety/Benchmarks (Samvelyan et al., 2024) 2024 Diverse adversarial prompt generation for automated safety stress testing.
Safety/Benchmarks (Wang et al., 2023a) 2023 Broad audit of trustworthiness, including safety and robustness, for GPT models.
Safety/Benchmarks (Tedeschi et al., 2024) 2024 Benchmarks red teaming effectiveness for assessing safety responses.
Fairness/Governance (Mittal et al., 2024) 2024 Dataset design practices aligned with fairness, privacy, and regulatory norms.
Fairness/Governance (Rodolfa et al., 2021) 2021 Minimal fairness–accuracy tradeoff in public policy ML tasks.
Fairness/Governance (Chouldechova, 2017) 2017 Disparate impact and fairness criteria in recidivism risk prediction.
Fairness/Governance (Barocas and Selbst, 2016) 2016 How big-data practices can create disparate impact and legal risk.
Fairness/Governance (Pessach and Shmueli, 2022) 2022 Survey of algorithmic fairness methods and evaluation across domains.
Fairness/Governance (Kaur et al., 2022b) 2022 Trustworthy AI principles and technical enablers for deployment.
Fairness/Governance (Dignum, 2019) 2019 Responsible AI framework grounded in ethics and governance.
Fairness/Governance (Laux et al., 2024) 2024 Trustworthiness concepts in relation to EU AI Act compliance.
Explainability/XAI (Ribeiro et al., 2016b) 2016 LIME: model-agnostic local explanations of predictions.
Explainability/XAI (Selvaraju et al., 2017) 2017 Grad-CAM for visual explanations in CNNs.
Explainability/XAI (Arrieta et al., 2020) 2020 Concepts, taxonomies, and challenges in explainable AI.
Explainability/XAI (Dwivedi et al., 2023) 2023 Review of XAI ideas/techniques with practical guidance.
Explainability/XAI (Saeed and Omlin, 2023) 2023 Meta-survey of XAI challenges and future directions.
Explainability/XAI (van der Waa et al., 2021) 2021 Rule-based vs. example-based explanations: fidelity and user value.
Explainability/XAI (Mosca et al., 2022) 2022 Review of SHAP-based explanation methods for NLP.
Explainability/XAI (Nauta et al., 2023) 2023 Systematizes quantitative evaluation protocols for explanations.
Privacy/Security (Dwork, 2006) 2006 Differential privacy as a formal framework for privacy protection.
Privacy/Security (McMahan et al., 2017) 2017 Federated learning for decentralized training with privacy benefits.
Privacy/Security (Nasr et al., 2023) 2023 Large-scale extraction of memorized training data from LMs.
Privacy/Security (Mattern et al., 2023) 2023 Membership inference via neighborhood comparison against LMs.
Privacy/Security (He et al., 2025a) 2025 Label-only membership inference attacks for pre-trained LMs.
Privacy/Security (Li et al., 2024c) 2024 Benchmark for privacy risks in LMs at multiple levels.
Privacy/Security (Mireshghallah et al., 2024) 2024 Privacy leakage tests using contextual-integrity evaluations.
Privacy/Security (Staab et al., 2024) 2024 Assesses large models for anonymization and de-identification tasks.
LLM Safety (Bai et al., 2022) 2022 Constitutional AI alignment to reduce harmful outputs.
LLM Safety (Weidinger et al., 2021) 2021 Maps ethical/social risks of LMs and mitigation avenues.
LLM Safety (Huang et al., 2024b) 2024 Positions trustworthiness dimensions and evaluation needs for LMs.
LLM Safety (Lin et al., 2024) 2024 Survey of debiasing and hallucination-reduction techniques.
LLM Safety (Wang et al., 2023c) 2023 ChatGPT applications, limitations, risks, and safeguards overview.